unicornquenn71:

Rewrite 5 2/3 as an improper fraction. 17/3 13/2 17/2 13/3

supie:

Do you know how to change to a improper fraction?

unicornquenn71:

i got 2

Angelwolf:

This is the process- Multiply the whole number by the denominator Add the answer from Step 1 to the numerator Write answer from Step 2 over the denominator

makaylab:

lol no

Downpour17:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Angelwolf This is the process- Multiply the whole number by the denominator Add the answer from Step 1 to the numerator Write answer from Step 2 over the denominator \(\color{#0cbb34}{\text{End of Quote}}\) Your very much correct this is the formula for solving a problem like this

unicornquenn71:

who ever doas this answer i will give them a medial

makaylab:

or you can times 5 x 2 + 3= and the bottom will be 3

makaylab:

5x3 +2 and the bottom will be 3

Angelwolf:

Step 1 Multiply the denominator by the whole number 2 × 5 = 10 Step 2 Add the answer from Step 1 to the numerator 10 + 3 = 13 Step 3 Write answer from Step 2 over the denominator 13/2

makaylab:

u do 5 x 3 +2 and the answer should look like this ?/3

unicornquenn71:

you got it right u get a medial angelwolf

Angelwolf:

Thx

Angelwolf:

Im right lol

makaylab:

was your answer 17/3

makaylab:

cuz mine sure was

makaylab:

before i even searched it up

Angelwolf:

the answer is 13/2

makaylab:

no it is 17/3 according to bing

Angelwolf:

Step 1 Multiply the denominator by the whole number 3 × 5 = 15 Step 2 Add the answer from Step 1 to the numerator 15 + 2 = 17 Step 3 Write answer from Step 2 over the denominator 17/3 = 17 3

makaylab:

17 3 =5 and 2 over 35 2 3 ≈5.667 Result17 over 3 equals 5 and 2 over 3 approximately 5.667 Step by step Step 1 of 1: Simplify. Simplify 5 and 2 over 35 2 3 = 17 over 3 17 3 Step 1 of 1: Simplify, sub-step a: Reduce fraction to lowest terms. Reduce fraction to lowest terms 1 is the greatest common divisor of 2 and 3. The result can't be further reduced. Step 1 of 1: Simplify, sub-step b: Convert mixed number to improper fraction. Convert mixed number to improper fraction 5 and 2 over 35 2 3 = ( 5 × 3 ) over 3 5 × 3 3 + 2 over 3 2 3 = ( 15 + 2 ) over 3 15 + 2 3 = 17 over 3 17 3

Angelwolf:

Omg! i switched the numerator and denominator im so sorry!!!

Angelwolf:

Thanks for the catch Makaylab

makaylab:

you are welcome that is all i am trying to say angel wolf

Angelwolf:

:)